# Degree Distribution based Spiking Graph Networks for Domain Adaptation

### v3

@@ Abstract
Spiking Graph Networks (SGNs) have garnered significant attraction from researchers and industry due to their ability to address energy consumption challenges in graph classification. However, SGNs are only effective for in-distribution data and cannot tackle out-of-distribution data. In this paper, we first propose the domain adaptation problem in SGNs, and introduce a novel framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA). The proposed DeSGDA addresses the spiking graph domain adaptation problem in three aspects: node degree-aware personalized spiking representation, adversarial feature distribution alignment, and pseudo-label distillation. First, we introduce the personalized spiking representation method for generating degree-dependent spiking signals. Specifically, the node degree determines the threshold of triggering a spike, allowing this personalized approach to capture more expressive information for classification. Then, we propose the graph feature distribution alignment module that is adversarially trained using membrane potential against a domain discriminator. Such an alignment module can efficiently maintain high performance and low energy consumption in the case of inconsistent distribution. Additionally, we extract consistent predictions across two spaces to create reliable pseudo-labels, effectively leveraging unlabeled data to enhance classification performance. Extensive experiments on benchmark datasets validate the superiority of the DeSGDA compared with competitive baselines.
@@ 1 Introduction
Spiking Graph Networks (SGNs) Zhu et al. (2022); Xu et al. (2021b) are a specialized artificial neural network engineered to process graph information by mimicking the human brain. SGNs transform static and real-valued graph features into discrete spikes by simulating neurons’ charging and discharging cycles, facilitating spike-based representations for graph node classification. Notably, SGNs excel in capturing semantic spiking representations with low energy consumption, which proves advantageous for event-based processing tasks Yao et al. (2021); Wang et al. (2024) such as object recognition Gu et al. (2020); Li et al. (2021b); Chen et al. (2024); Yin et al. (2024a), real-time data analysis Zhu et al. (2020); Bauer et al. (2019); Yin et al. (2023b), and graph classification Li et al. (2023); Zhu et al. (2022); Liu et al. (2024c); Yin et al. (2022a).
Currently, SGNs are usually tested within the same distribution as the training dataset Li et al. (2023); Yin et al. (2024a); Duan et al. (2024). However, in realistic scenarios, the testing set can have different distributions from the training set, and such a distribution shift may degrade performance. For instance, Electroencephalography (EEG) data Binnie and Prior (1994); Biasiucci et al. (2019), typically represented as a graph structure with nodes for neurons and edges for connections, is ideally processed by bio-inspired SGNs that mimic neuronal charging and discharging. Despite the suitability, EEGs often exhibit varying distributions over time or among different groups Zhao et al. (2020, 2021); Wang et al. (2022), leading to suboptimal performance of models trained on specific distributions when applied to others. This significant issue underscores the necessity of exploring domain adaptation for spiking graphs. Traditionally, SNNs transfer learning methods Zhan et al. (2021); Yin et al. (2023c, 2024b); Pang et al. (2023) have been applied in event-based or computer vision scenarios. However, there’s no existing research on spiking graph domain adaptation.
In this paper, we address the development of energy-efficient SGNs tailored for scenarios involving distribution shifts. Both domain adaptation and SGNs are particularly well-suited for real-world applications where data distributions vary across environments, and efficient processing of graph-structured, dynamic data under resource constraints is crucial. These challenges are common across numerous fields that require solutions capable of handling distribution shifts while minimizing energy consumption. However, designing an effective spiking graph domain adaptation framework is non-trivial due to the following major challenges: (1) How to meticulously design an SGN under the circumstance of domain shift? SGNs usually utilize a global threshold for the firing of each node Xu et al. (2021a); Yin et al. (2024a); Zhao et al. (2024). However, we observe that the degree of each node influences the difficulty of triggering spikes. Specifically, nodes with high degrees can integrate more information from neighbors, making it easier for membrane potential to accumulate and trigger a spike. Conversely, nodes with lower degrees are more challenging to reach the firing threshold, denoted as the inflexible architecture challenge. (2) How to design a framework that effectively addresses spiking graph domain adaptation for classification? Current research primarily focuses on graph node classification within the same distribution Li et al. (2023); Yao et al. (2023a); Duan et al. (2024). However, spike-based graph classification under domain shift remains unexplored. (3) How to guarantee the stability of the proposed framework? Though some works have been proposed to address the spiking transfer learning challenges Zhan et al. (2021); Zhang et al. (2021); Zhan et al. (2024), there is still no theoretical research on spiking graphs under domain shift.
To tackle these challenges, we propose a framework named Degree-aware Spiking Graph Domain Adaptation for Classification (DeSGDA), which comprises three components: degree-aware personalized spiking representation, graph feature distribution alignment, and pseudo-label distillation. To address the first challenge, we establish variable node thresholds based on their degrees. By adaptively updating these thresholds, we achieve a more expressive and personalized spiking representation for each node. Then, we introduce an adversarial feature distribution alignment module trained using membrane potential against a domain discriminator. To further enhance performance, we extract consistent predictions from different spaces to generate reliable pseudo-labels. Additionally, to explore the generalization ability of the proposed DeSGDA, we propose the error bound for spiking graph domain adaptation and demonstrate that our pseudo-label distillation module effectively reduces this upper bound. In summary, we utilize simple yet effective techniques to address a novel problem while providing an insightful analysis of the background mechanisms and capabilities of our proposed method.
Our contributions can be summarized as follows:
  • Problem Formulation: We first introduce the problem of spiking graph domain adaptation for classification, which is non-trivial due to the challenges of the inflexible architecture of SGNs and theoretical deficiency.
  • Novel Architecture: We propose DeSGDA, a framework that efficiently learns personalized spiking representations for nodes using degree-aware thresholds and aligns domain distributions through adversarial training on membrane potential. Furthermore, we utilize pseudo-label distillation to improve the performance further.
  • Theoretical Analysis: To guarantee the stability of DeSGDA, we provide theoretical proof of the error bound for spiking graph domain adaptation. Furthermore, we demonstrate that DeSGDA maintains a lower theoretical bound than standard spiking graph domain adaptation through the use of the pseudo-label distillation module.
  • Extensive Experiments. We evaluate the proposed DeSGDA on extensive spiking graph domain adaptation learning datasets, which shows that our proposed DeSGDA outperforms the variety of state-of-the-art methods.
@@ 2 Related work
Figure 1: An overview of the proposed DeSGDA. To achieve personalized spiking representations, DeSGDA employs adaptive thresholds based on node degrees, enabling the generation of tailored spiking representations. To align domain distributions, DeSGDA leverages adversarial training on membrane potentials to counter domain discrimination. Furthermore, DeSGDA utilizes pseudo-labeling to identify and select reliable samples, thereby enhancing overall model performance.
Spiking Graph Networks (SGNs). SGNs are a specialized type of neural network that combines Spiking Neural Networks (SNNs) with Graph Neural Networks (GNNs), preserving energy efficiency while achieving competitive performance in various graph tasks Li et al. (2023); Yao et al. (2023a, b); Duan et al. (2024); Liang et al. (2023). Existing research on SGNs focuses on capturing the dynamic temporal information within graphs and enhancing scalability. For instance, Xu et al. (2021a) utilizes spatial-temporal feature normalization within SNNs to effectively process dynamic graph data, ensuring robust learning and improved predictive performance. Zhao et al. (2024) propose a method that dynamically adapts to evolving graph structures and relationships through a novel architecture that updates node representations in real time. Additionally, Yin et al. (2024a) adapts SNNs to dynamic graph settings and employs implicit differentiation for the node classification task. However, existing methods still suffer from data distribution shift issues when training and testing data come from different domains, resulting in degraded performance and generalization. To address this, we propose a novel domain adaptation method based on SGNs to tackle these challenges.
Spiking Transfer Learning. Spiking transfer learning focuses on adjusting SNNs to handle data distribution shifts across various domains more effectively. Recent advances in spiking transfer learning have been extensively applied in vision tasks, significantly enhancing model performance while maintaining energy efficiency Zhan et al. (2021); Zhang et al. (2021); Zhan et al. (2024). For instance, Guo et al. (2024) leverages a Jaccard attention mechanism within SNNs to adapt to target domains without requiring source domain data. Similarly, He et al. (2024) facilitates the transfer of learned representations from static to dynamic event-based domains by adapting SNNs to process temporal information. Additionally, Zhan et al. (2024) converts RGB images into spike-based neuromorphic data, enabling SNNs to process visual information across various domains efficiently. However, the inherent difficulty of complex graph topologies makes it infeasible to apply spiking transfer learning to SGNs directly. To this end, we introduce a specialized domain adaptation method specifically tailored for SGNs.
@@ 3 Preliminaries
Bound for Graph Domain Adaptation (GDA). Applying GDA with optimal transport (OT), if the covariate shift holds on representations that $\mathbb{P}_{S}(Y|Z)=\mathbb{P}_{T}(Y|Z)$, the target risk $\epsilon_{T}(h,\hat{h})$ is bounded with the theorem:
@@
You et al. (2023) Assuming that the learned discriminator is $C_{g}$-Lipschitz continuous as described in Redko et al. (2017), and the graph feature extractor $f$ (also referred to as GNN) is $C_{f}$-Lipschitz that $||f||_{Lip}=\max_{G_{1},G_{2}}\frac{||f(G_{1})-f(G_{2})||_{2}}{\eta(G_{1},G_{2})}=C_{f}$ for some graph distance measure $\eta$. Let :=$\mathcal{H}:=\{h:\mathcal{G}\rightarrow\mathcal{Y}\}$ be the set of bounded real-valued functions with the pseudo-dimension $Pdim(\mathcal{H})=d$ that $h=g\circ f\in\mathcal{H}$, with probability at least $1-\delta$ the following inequality holds:
@@
$\epsilon_{T}(h,\hat{h})\leq$ $\ \hat{\epsilon}_{S}(h,\hat{h})+\sqrt{\frac{4d}{N_{S}}\log\left(\frac{eN_{S}}{d}\right)+\frac{1}{N_{S}}\log\left(\frac{1}{\delta}\right)}$
$+2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))+\omega,$
where the (empirical) source and target risks are $\hat{\epsilon}_{S}(h,\hat{h})=\frac{1}{N_{S}}\sum_{n=1}^{N_{S}}|h(G_{n})-\hat{h}(G_{n})|$ and $\epsilon_{T}(h,\hat{h})=\mathbb{E}_{\mathbb{P}_{T}(G}\{|h(G)-\hat{h}(G)|\}$, respectively, where $\hat{h}:\mathcal{G}\rightarrow\mathcal{Y}$ is the labeling function for graphs and $\omega=\min_{||g||_{Lip}\leq C_{g},||f||_{Lip}\leq C_{f}}\{\epsilon_{S}(h,\hat{h})+\epsilon_{T}(h,\hat{h})\}$. The first Wasserstein distance is defined as Villani et al. (2009): $W_{1}(\mathbb{P},\mathbb{Q})=\sup_{||g||_{Lip}\leq 1}\left\{\mathbb{E}_{\mathbb{P}_{S}(Z)}g(Z)-\mathbb{E}_{\mathbb{P}_{T}(Z)}g(Z)\right\}$.
The comprehensive justification of the OT-based graph domain adaptation bound demonstrates that the generalization gap relies on both the domain divergence $2C_{f}C_{g}W_{1}(\mathbb{P}_{S}(G),\mathbb{P}_{T}(G))$ and model discriminability $\omega$.
Spiking Graph Networks. In contrast to traditional artificial neural networks, SGNs Xu et al. (2021a); Zhu et al. (2022) convert input data into binary spikes over time, with each neuron in the SGNs maintaining a membrane potential that accumulates input spikes. A spike is produced as an output when the membrane potential exceeds a threshold, which is formulated as:
$u_{\tau+1,i}=\lambda(u_{\tau,i}-V_{th}s_{\tau,i})+\sum_{j}w_{ij}\mathcal{A}\left(A,s_{\tau,j}\right)+b,\vspace{-4pt}$
$s_{\tau+1,i}=\mathbb{H}(u_{\tau+1,i}-V_{th}),$ (1)
where $\mathbb{H}(x)$ is the Heaviside function, which is the non-differentiable spiking function, $\mathcal{A}$ is the graph aggregation operation, and $A$ is the adjacency matrix of a graph. $s_{\tau,i}$ denotes the binary spike train of neuron $i$, and $\lambda$ is the constant. $w_{ij}$ and $b$ are the weights and bias of each neuron.
4 Methodology

### v4

# Degree-Conscious Spiking Graph for Cross-Domain Adaptation

@@ Abstract
Spiking Graph Networks (SGNs) have demonstrated significant potential in graph classification by emulating brain-inspired neural dynamics to achieve energy-efficient computation. However, existing SGNs are generally constrained to in-distribution scenarios and struggle with distribution shifts. In this paper, we first propose the domain adaptation problem in SGNs, and introduce a novel framework named Degree-Consicious Spiking Graph for Cross-Domain Adaptation (DeSGraDA). DeSGraDA enhances generalization across domains with three key components. First, we introduce the degree-conscious spiking representation module by adapting spike thresholds based on node degrees, enabling more expressive and structure-aware signal encoding. Then, we perform temporal distribution alignment by adversarially matching membrane potentials between domains, ensuring effective performance under domain shift while preserving energy efficiency. Additionally, we extract consistent predictions across two spaces to create reliable pseudo-labels, effectively leveraging unlabeled data to enhance graph classification performance. Furthermore, we establish the first generalization bound for SGDA, providing theoretical insights into its adaptation performance. Extensive experiments on benchmark datasets validate that DeSGraDA consistently outperforms state-of-the-art methods in both classification accuracy and energy efficiency.
@@ 1 Introduction
Spiking Graph Networks (SGNs) (Zhu et al., 2022; Xu et al., 2021b) as a specialized neural network combining Spiking Neural Networks (SNNs) (Gerstner and Kistler, 2002; Maass, 1997) with Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Scarselli et al., 2009; Chen et al., 2024), have emerged as a groundbreaking paradigm in artificial neural networks, uniquely designed to process graph-structured data by mimicking the bio-inspired mechanisms of the human brain. SGNs convert graph features into binary spiking signals, replacing matrix multiplications with simple additions to achieve high energy efficiency. They further exploit temporal spiking representations, encoding information in spike timing to enable asynchronous, event-driven processing. It is particularly critical for applications where energy consumption is a bottleneck, such as real-time brain-computer interfaces (Kumar et al., 2022; Nason et al., 2020), large-scale sensor networks (Yao et al., 2021; Wilson et al., 2024), and temporal analysis (Yin et al., 2024; Zhu et al., 2024; Zhou et al., 2021).
Despite their potential for energy-efficient graph representation, existing SGNs are primarily studied under closed-world assumptions, where source and target data share identical distributions (Li et al., 2023; Yin et al., 2024; Duan et al., 2024). This assumption is inadequate for many real-world scenarios, such as cross-subject brain–computer interfaces (BCIs) (Binnie and Prior, 1994; Biasiucci et al., 2019) where distribution shifts can significantly degrade performance(Zhao et al., 2020, 2021; Wang et al., 2022). Although recent advancements in transfer learning for SNNs have shown promise in vision tasks by leveraging neuromorphic adaptations (Zhang et al., 2021; Zhan et al., 2024; Guo et al., 2024), they are primarily designed for grid-like inputs and fail to generalize to graph data. The non-Euclidean nature and inherent irregularity of graphs introduce fundamental challenges (Bronstein et al., 2017; Wang et al., 2024b; Yao et al., 2023b), as existing methods neglect the topological dependencies and message-passing mechanisms crucial for effective graph learning. Consequently, directly applying these methods to SGNs results in suboptimal adaptation under domain shifts.
In this paper, we investigate the development of energy-efficient SGNs for scenarios involving distribution shifts. However, designing an effective domain adaptation framework for SGNs poses several fundamental challenges: (1) How to adapt spiking representations to account for the structural diversity of graph-structured data? Traditional SGNs typically assign a fixed firing threshold to all nodes, ignoring the structural diversity in graphs (Xu et al., 2021a; Yin et al., 2024). This uniform treatment leads to under-activation in nodes with fewer connections, where important features are missed, and over-activation in highly connected nodes, where excessive signals distort the learned representation. Both cases reduce the model’s representational effectiveness. (2) How to design domain adaptation strategies that account for the temporal-based representations? Unlike static neural models, SGNs encode information through temporal spike sequences, making them more sensitive to domain shifts (Zhou et al., 2023; Zhan et al., 2021). Existing methods fail to address how these shifts impact the spike sequences, resulting in suboptimal alignment across domains. (3) How to theoretically characterize and bound the generalization error of SGNs under domain shift? Despite empirical advances (Zhang et al., 2021; Zhan et al., 2024), the theoretical understanding of SGN domain adaptation remains limited. Without a principled framework to quantify generalization under distribution shift, it is difficult to design adaptation methods with guaranteed performance.
To tackle these challenges, we propose a framework called Degree-Conscious Spiking Graph for Cross-Domain Adaptation (DeSGraDA). This has three components: (i) degree-conscious spiking representation, which assigns variable firing thresholds to nodes based on their degrees, enabling adaptive control over spiking sensitivity. This degree-conscious mechanism balances the firing frequencies of high- and low-degree nodes, preventing information loss in sparsely connected nodes and avoiding distortion from excessive signals in highly connected ones, thus enhancing the expressiveness of the spiking representations; (ii) temporal distribution alignment, which explicitly aligns the time-evolving spiking representations between the source and target domains. By leveraging membrane potential dynamics as evolving signals, the model captures domain-specific patterns, improving robustness to temporal shifts; and (iii) pseudo-label distillation assigns reliable pseudo-labels by aligning consistent predictions from shallow and deep network layers. We also demonstrate that this pseudo-label distillation module can effectively reduce a generalization bound tailored for spiking graph domain adaptation. In summary, DeSGraDA provides a simple yet effective solution to a novel and underexplored problem, and offers deep insights into the spiking graph domain adaptation.
Our contributions can be summarized as follows: (1) Problem Formulation: We first introduce the spiking graph domain adaptation for graph classification, highlighting the challenges posed by the inflexible threshold mechanism of SGNs and theoretical limitations that hinder effective adaptation. (2) Novel Architecture and Theoretical Analysis: We propose DeSGraDA, a framework combining degree-conscious spiking representation and temporal distribution alignment. Moreover, we provide a generalization bound for spiking graph domain adaptation. (3) Extensive Experiments. We evaluate the proposed DeSGraDA on extensive spiking graph domain adaptation learning datasets, demonstrating that it can outperform various state-of-the-art methods.
@@ 2 Related Work
Domain Adaptation (DA). DA transfers knowledge from a labeled source domain to an unlabeled target domain by mitigating the distributional shift between the two domains (Redko et al., 2017; Long et al., 2018; Shen et al., 2018). It has been widely applied to vision and language tasks (Wei et al., 2021b, a; Shi et al., 2024). Recently, DA has been extended to graph data to address the unique challenges posed by complex relationships, leading to the emergence of Graph Domain Adaptation (GDA) (Yin et al., 2023; Liu et al., 2024a; Cai et al., 2024). Most existing GDA approaches first leverage GNNs to generate node and graph representations (Zhu et al., 2021; Yin et al., 2022; Liu et al., 2024b), followed by adversarial learning to implicitly align feature distributions and reduce domain discrepancies. They also apply structure-aware strategies to explicitly align graph-level semantics and topological structures, thereby enhancing generalization across diverse graph domains (You et al., 2023; Liu et al., 2024b; Luo et al., 2024). However, GDA remains under-explored in the context of spiking graphs, where energy efficiency becomes a critical requirement for real-world applications. To bridge this gap, we introduce the novel problem of Spiking Graph Domain Adaptation (SGDA), extending GDA to spiking graphs for energy-efficient domain adaptation.
Spiking Graph Networks (SGNs). SGNs are a specialized neural network combining SNNs (Gerstner and Kistler, 2002; Maass, 1997) with GNNs, preserving energy efficiency while achieving competitive performance on various tasks (Li et al., 2023; Yao et al., 2023a; Duan et al., 2024). Existing research on SGNs focuses on capturing temporal information within graphs and enhancing scalability. For instance, Xu et al. (2021a) utilizes spatial-temporal feature normalization within SNNs to effectively process dynamic graph data, ensuring robust learning and improved performance. Zhao et al. (2024) proposes a method that adapts to evolving graph structures through a novel architecture that updates node representations in real time. Additionally, Yin et al. (2024) adapts SNNs to dynamic graph settings and employs implicit differentiation for the node classification task. However, existing methods still suffer from data distribution shift issues when the training and test data come from different domains, resulting in degraded performance and generalization. To tackle these challenges, we propose a novel domain adaptation method for spiking graph networks.
@@ 3 Preliminaries
Problem Setup. Given a graph $G=(V,E,\mathbf{X})$ with node set $V$, edge set $E$, and node attribute matrix $\mathbf{X}$. To construct spiking-compatible inputs for SGNs, we sample the binary features $S$ from the Bernoulli distribution with probability of $\mathbf{X}$ (i.e., $S\sim Ber(\mathbf{X})$) as input of SGNs (Zhao et al., 2024). In this paper, we focus on the problem of spiking GDA for graph classification. The source domain $\mathcal{D}^{s}=\{(G_{i}^{s},y_{i}^{s})\}_{i=1}^{N_{s}}$ is labeled, where $N^{s}$ is the number of source-domain graphs $\{G_{i}^{s}\}$ and $y_{i}^{s}$ is the label of $G_{i}^{s}$. The target domain $\mathcal{D}^{t}=\{G_{j}^{t}\}_{j=1}^{N_{t}}$ is unlabeled and contains $N^{t}$ graphs. Both domains share the same label space $\mathcal{Y}$ but can have different graph topologies or attribute distributions.
Domain Adaptation with Optimal Transport (OT). Following You et al. (2023), we factorize a trained model $h$ as $g\circ f$, where $f:\mathcal{D}\mapsto\mathbb{R}^{d}$ is the feature extractor ($Z=f(\mathcal{D})$) and $g:\mathbb{R}^{d}\mapsto\mathcal{Y}$ is the discriminator ($Y=g(Z)$). For simplicity, we focus on binary classification where $\mathcal{Y}=[0,1]$. Denote the classifier that predicts labels from the feature representation as $\hat{g}:\mathbb{R}^{d}\mapsto\mathcal{Y}$. The (empirical) source and target risks are given by $\hat{\epsilon}_{S}(g,\hat{g})=\frac{1}{N_{s}}\sum_{n=1}^{N_{s}}|g(Z_{n})-\hat{g}(Z_{n})|$ and $\epsilon_{T}(g,\hat{g})=\mathbb{E}_{\mathbb{P}_{T}(Z)}|g(Z)-\hat{g}(Z)|$, respectively. DA with optimal transport (OT) (Villani et al., 2008) addresses covariate shift by finding the optimal way to transport masses between the two source and target distributions while minimizing the transport cost. The following Theorem shows that the generalization gap depends on both the domain divergence $2C_{g}W_{1}(\mathbb{P}_{S}(Z),\mathbb{P}_{T}(Z))$ and model discriminability $\omega$.
@@
(Redko et al., 2017; Shen et al., 2018) Assume that the discriminator $g$ is $C_{g}$-Lipschitz. [^1] The Lipschitz norm $\|g\|_{Lip}=\max_{Z_{1},Z_{2}}\frac{|g(Z_{1})-g(Z_{2})|}{\rho(Z_{1},Z_{2})}=C_{g}$ for some given distance function $\rho$. Let :=$\mathcal{H}:=\{g:\mathcal{Z}\rightarrow\mathcal{Y}\}$ (where $\mathcal{Z}$ is the feature space) be the set of bounded real-valued functions with pseudo-dimension $Pdim(\mathcal{H})=d$. For any $g\in\mathcal{H}$, the following holds with probability at least $1-\delta$:
@@
$\epsilon_{T}(g,\hat{g})\leq$ $\hat{\epsilon}_{S}(g,\hat{g})+\sqrt{\frac{4d}{N_{S}}\log\left(\frac{eN_{S}}{d}\right)+\frac{1}{N_{S}}\log\left(\frac{1}{\delta}\right)}+2C_{g}W_{1}(\mathbb{P}_{S}(Z),\mathbb{P}_{T}(Z))+\omega,$
where $\omega=\min_{\|g\|_{Lip}\leq C_{g}}\{\epsilon_{S}(g,\hat{g})+\epsilon_{T}(g,\hat{g})\}$ is the discriminative ability to capture source and target data, and $W_{1}(\mathbb{P},\mathbb{Q})$ is the distribution divergence defined in (Redko et al., 2017; Shen et al., 2018).
Applying OT-based DA methods to graph data introduces additional challenges due to the non-Euclidean nature of graphs and intricate dependencies between nodes. You et al. (2023) extends the OT-based DA framework to graphs, and provides a generalization bound for GDA. Details are in Appendix A.
Spiking Graph Networks (SGNs). Spiking Neural Networks (SNNs) are brain-inspired models that communicate through discrete spikes rather than continuous-valued activations (Maass, 1997; Gerstner and Kistler, 2002; Bohte et al., 2000). This provides significant advantages in temporal information processing and energy efficiency. More details are in Appendix B. SGNs are a specialized form of SNNs tailored for graph data (Xu et al., 2021a; Zhu et al., 2022), where each node is modeled as a spiking neuron. The membrane of each node evolves based on both the temporal spiking dynamics and the graph’s structural connectivity. Let $u_{\tau,i}$ be the membrane potential of node $i$ at latency step $\tau$. SGNs first updates the membrane potentials via input aggregation:
$u_{\tau+1,i}=\lambda(u_{\tau,i}-V_{th}s_{\tau,i})+\sum\nolimits_{j}w_{ij}\mathcal{A}(A,s_{\tau,j})+b_{i},$ (1)
where $s_{\tau,i}$ is the spiking representation, $\lambda\in(0,1)$ is the leak factor, $V_{th}$ is the firing threshold, $b_{i}$ is the bias, $w_{ij}$ is the synaptic weight from node $j$ to node $i$, $A$ is the adjacency matrix, and $\mathcal{A}$ is the graph-based aggregation operation. Next, spikes are triggered through thresholding:
$s_{\tau+1,i}=\mathbb{H}(u_{\tau+1,i}-V_{th}),\quad u_{\tau+1,i}=(1-s_{\tau+1,i})u_{\tau+1,i}+s_{\tau+1,i}V_{reset},$ (2)
where $V_{\text{reset }}$ is the reset potential, and $\mathbb{H}(x)$ is the Heaviside step function which serves as the non-differentiable spiking function. Finally, the neuron is reset upon firing.
4 Proposed Methodology
Figure 1: Overview of the proposed DeSGraDA. Degree-Conscious Spiking Representation generates source and target domain spiking representations by adapting neuron firing threshold based on node degrees. To align domain distributions, Temporal Distribution Alignment leverages adversarial training on temporal membrane potential to counter domain discrimination. Furthermore, Pseudo-labe Distillation is employed to identify reliable target samples and reinforce overall model performance.